OpenStudy (anonymous):

a woman can do a certain job in 10 hours and her younger daughter can do it in 12 hours. after the woman and her younger daughter have been working for 1 hour, they are join by the older daughter and the three complete the job in 3 more hours. how will it take the older daughter to do the jib by herself. please help i'm really confused.

OpenStudy (anonymous):

Let'a divide the job into 60 parts (LCM of 10 and 12). The woman can do 6 parts an hour. The younger daughter can do 5 parts an hour. Together, they do 11 parts an hour. 49 parts left when the older daughter joins. Let the older daughter's work speed be Y. So 3*(11+Y)=49 33+3Y=49 3Y=49-33 3Y=16 The older daughter does 5 and a third parts an hour. So 60/5.333...=11.25 hours

OpenStudy (anonymous):

thank you!